OpenStudy (anonymous):

Please help me!!!! solve the system using elimination x+ y + +z =1 2x-y+2x=-5 -x+2y-z = 4

hero (hero):

[1 1 1 1] [2 -1 2 -5] [-1 2 -1 4] Set up matrix

OpenStudy (anonymous):

i have to use elimination

hero (hero):

Good luck with that

OpenStudy (anonymous):

i am so confused on how to do these can someone please help!!!

OpenStudy (anonymous):

x+ y +z =1 2x-y+2z=-5 -x+2y-z = 4 x + y+z=1 -x+2y-z=4 3y = 5 2x-y+2z=-5 (-x+2y-z=4)2 -2x -4y-2z = 8 -5y = 3 3y = 5 -5y = 3 -2y = 8 y = -4 ^^^that's your the y part of the ordered triplet, now you just need to find the x and z values

OpenStudy (anonymous):

thanks sooooooooo...... much

OpenStudy (anonymous):

could you help me with one more?

OpenStudy (anonymous):

x+ y + +z =1 2x-y+2x=-5 -x+2y-z = 4 x+ y +z =1 2x-y+2z=-5 3x+3z =-4 (2x-y+2z=-5)2 4x-2y+4z=-10 -x+2y-z = 4 3x+3z=-6 3x+3z=-4 (3x+3z=-6)-1 -3x-3z=6 0 does not equal 2, so there is no solution for the problem